OpenStudy (anonymous):

Solve: 7x - 2(x + 1) = 6x + 14 16 12 -16 -12

jhonyy9 (jhonyy9):

7x-2x-2=6x+14 5x-2=6x+14 -16=x

jhonyy9 (jhonyy9):

moneybird there is -2 not -1

OpenStudy (anonymous):

7x - 2(x + 1) = 6x + 14 7x -2x - 2 = 6x + 14 5x-2 = 6x + 14 -16 = x i got the same as jhonyy9

jhonyy9 (jhonyy9):

yes moyo this is right sure

OpenStudy (anonymous):

johnyy9's answer is correct. moneybird made a mistake, because he didn't multiplied 2 by 1 in "-2(x+1)".